OpenStudy (anonymous):

a>b>c>0, what could be the possible values of a,b,c if abc=ab+bc+ca

OpenStudy (jlg030597):

welcome to Open Study.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

one set of values are 3,6,10

ganeshie8 (ganeshie8):

3,6,10 dont satisfy the given equation right ?

ganeshie8 (ganeshie8):

And, any restrictions given on a,b,c ? are they integers ?

OpenStudy (anonymous):

no more info given....sorry the given set is wrong, 3,6.10 :(

ganeshie8 (ganeshie8):

i think there wil be finite # of solutions only if a,b,c are integers

OpenStudy (anonymous):

a,b,c cant be fractions, 'cos then i divided by the number exceeds 1 in that case

OpenStudy (anonymous):

can you think of a way to get the solutions?

ganeshie8 (ganeshie8):

\(abc=ab+bc+ca\) divide abc thru out \(1=\dfrac{1}{a} + \dfrac{1}{b} + \dfrac{1}{c}\) only one solution : \((2, 3, 6)\)

ganeshie8 (ganeshie8):

no more solutions possible in integers

OpenStudy (anonymous):

how did you make sure?

ganeshie8 (ganeshie8):

Notice that \(\dfrac{1}{3} +\dfrac{1}{3}+ \dfrac{1}{3} = 1\), mess with it - fix one value and see what values other two can take

ganeshie8 (ganeshie8):

here is a nice explanation : http://krismanohar.com/blog/?p=157

OpenStudy (anonymous):

Thank you for all the help :)

ganeshie8 (ganeshie8):

you're welcome :)

OpenStudy (anonymous):

i wish there were a generalized solution for solving for a1, a2, a3...an...with the condition 1/a1+1/a2 ....=1 like mentioned in the blog you posted

ganeshie8 (ganeshie8):

yeah should be possible... definitely worth trying xD

ganeshie8 (ganeshie8):

solve in distinct positive integers : \(\dfrac{1}{a_1} + \dfrac{1}{a_2} + ... + \dfrac{1}{a_n} = 1\)

OpenStudy (anonymous):

trying with 4 integers...

ganeshie8 (ganeshie8):

try (2,4,6,12)

ganeshie8 (ganeshie8):

it looks extremely difficult to cookup a formula

OpenStudy (anonymous):

wow....u got dat

OpenStudy (anonymous):

it feels frightening to just plug and play with numbers

OpenStudy (anonymous):

Hey Ganeshie8, found a nice blog on how to solve such ones, http://mathforum.org/library/drmath/view/56821.html

ganeshie8 (ganeshie8):

Ahh nice :) so basically they're doing : \[ 1 = \dfrac{1}{2} + \left(1 - \dfrac{1}{2}\right)\]

ganeshie8 (ganeshie8):

\[ 1 = \dfrac{1}{2} + \dfrac{1}{3} + \left(1 - \left[\dfrac{1}{2} + \dfrac{1}{3}\right]\right)\]

ganeshie8 (ganeshie8):

\[ 1 = \dfrac{1}{2} + \dfrac{1}{3} + \dfrac{1}{7} + \left(1 - \left[\dfrac{1}{2} + \dfrac{1}{3}+ \dfrac{1}{7}\right]\right)\]

ganeshie8 (ganeshie8):

^gives the solution : \((2,3,7, 42)\)

OpenStudy (anonymous):

yes!!! Thank God i found that site...a blessing!!! Thank you also for being there, supporting!!!

ganeshie8 (ganeshie8):

np :) lol i was almost giving up... and u gave me that site xD thank you too :))

OpenStudy (anonymous):

for 5 ints you have: (2,3,7,43,1806)

ganeshie8 (ganeshie8):

yeah and a plenty more http://www.wolframalpha.com/input/?i=1+-%281%2F2%2B1%2F3%2B1%2F7%2B1%2F43%29

OpenStudy (anonymous):

That's nice site, thank you again!

ganeshie8 (ganeshie8):

i think 1806 is max value possible for a/b/c/d/e, so it looks like there will not be infinite combinations for 5 ints

ganeshie8 (ganeshie8):

for 4 ints, also i think 42 is the max possible value for a/b/c/d, so these solutions are also finite

ganeshie8 (ganeshie8):

but not fully sure... il need to think bit more thoroughly

OpenStudy (anonymous):

by the way, according to the method given on Dr.Math for 5 ints you have: (2,3,7,43,1806) should work...not sure why the difference ends up in 1/1806...oh...that's b'cos it takes the decimals and computes the answer so there's a difference coming of 1/1806

OpenStudy (anonymous):

the above post was wrong...i messed up thinking 1/1806 is got after unit factors of 5 ints are subtracted from 1..

OpenStudy (anonymous):

but Dr.Math said that there are infinite solutions possible?

ganeshie8 (ganeshie8):

yeah need to think a bit more - consider 4ints : can there be solutions such that any one of a,b,c,d is greater than 42 ?

ganeshie8 (ganeshie8):

im thinking the value 42 is the upper limit for 4ints, il think more on this later... for now i need to go work on something else...

OpenStudy (anonymous):

sure, TC

ganeshie8 (ganeshie8):

solution : \((a, b, c, d)\) all \(\le 42\) is what my best guess is. see if u can give a counter example

ganeshie8 (ganeshie8):

have fun :)